Given a mean of 8 and a standard deviation of 0.7, what is the z-score of the value 9 rounded to the nearest tenth?
Answer by jim_thompson5910(35256)

x = 9
mu = 8
sigma = 0.7


z = (x-mu)/sigma

z = (9-8)/0.7

z = 1/0.7

z = 1.42857142857142

z = 1.4